Normal saline 1 liter to run from 0800-1800 o’clock . The patient went down for a procedure at 1400 o’clock and the IV had to be disconnected. The patient came back at 1600 o’clock. The doctor told you to complete the infusion at 1800 o’clock.
1-what is the initial flow rate?
2-recalculalte the flow rate
Answer #1

1)Given

Normal saline:1 litre (1000ml)

Start time: 8 am

End time : 6pm

Total time: 10 hours

Stopped between 2pm to 4pm ( 2 hours)

1) flow rate:Total volume*15/10*60

1000*15/10*60

=25 drops /minute is the flow rate till 6pm

so till 2pm the amount transfused is

25=Total volume*15/6*60

25=TV/6*4

25=TV/24

TV=25*24

Total volume infused is 600 ml at the Raye of 25 drips per minute till 2 pm

2)Balance 400 ml

So after 4 pm the flow rate will be

flow rate= 400*15/2*60

=400/8

=50 drops/ minute for 2 hours from 4 pm to 6pm